### v2

Abstract
Self-supervised deep learning has accelerated 2D natural image analysis but remains difficult to translate into 3D MRI, where data are scarce and pre-trained 2D backbones cannot capture volumetric context. We present a sequence-invariant self-supervised framework leveraging quantitative MRI (qMRI). By simulating multiple MRI contrasts from a single 3D qMRI scan and enforcing consistent representations across these contrasts, we learn anatomy-centric rather than sequence-specific features. This yields a robust 3D encoder that performs strongly across varied tasks and protocols. Experiments on healthy brain segmentation (IXI), stroke lesion segmentation (ARC), and MRI denoising show significant gains over baseline SSL approaches, especially in low-data settings (up to +8.3% Dice, +4.2 dB PSNR). Our model also generalises effectively to unseen sites, demonstrating potential for more scalable and clinically reliable volumetric analysis. All code and trained models are publicly available.
editors: Under Review for MIDL 2025
1 Introduction
Deep learning has advanced medical image analysis across a range of tasks, from image registration [Balakrishnan et al. (2019)] to semantic segmentation [Chalcroft et al. (2023)] and visual question-answering [Narayanan et al. (2024)]. However, unique challenges arise when working with 3D MRI data, including increased computational demands and the difficulty of applying 2D pre-trained models to volumetric contexts [Ma et al. (2024)]. Although large-scale 3D datasets [Li et al. (2024)] and models [Wu et al. (2024)] have recently emerged, fine-tuning them for specific clinical tasks remains non-trivial due to inevitable domain shifts [Yang et al. (2024)].
Self-supervised learning (SSL) offers a promising means of learning robust representations without the need for large labelled datasets [Chaitanya et al. (2020), Huang et al. (2023)]. Yet, existing SSL methods often treat each MRI sequence as a separate domain, neglecting the shared anatomical information across contrast variations. In contrast, we leverage the observation that different MRI sequences, despite their unique contrast properties, encode the same underlying anatomy. By incorporating physics-based knowledge of MRI signal generation into an SSL framework, we can learn representations that are robust to sequence changes while preserving anatomical fidelity.
In this paper, we make the following contributions:
  • We propose a sequence-invariant self-supervised learning framework based on physics-driven contrast simulation, enabling robust 3D MRI representations across multiple sequences.
  • We show significant performance gains in low-data regimes, with up to 8.3% Dice improvement in segmentation and a 4.2 dB PSNR increase for denoising when using just 1% of the training data.
  • We demonstrate enhanced cross-sequence generalisation and site adaptation from a single pre-trained encoder, particularly in out-of-domain scenarios.
  • We provide comprehensive evaluations on three diverse tasks - healthy brain segmentation, stroke lesion segmentation, and image denoising - highlighting the clinical utility of our approach.
Our method addresses key problems in medical imaging by enabling robust feature learning across different sites and sequences, even with limited annotated data. This work takes a step towards developing more generalisable and clinically applicable models. We release all code and backbone weights at github.com/liamchalcroft/contrast-squared.
2 Related Work
We briefly review three core areas that underpin this work: contrastive learning, robust representations in 3D medical imaging, and quantitative MRI (qMRI).
2.1 Contrastive Learning
Self-supervised learning (SSL) can leverage unlabelled data by creating proxy tasks that encourage useful invariances in learned representations. Techniques include predictive coding (van den Oord et al., 2019; Hénaff et al., 2020), masked image modelling (He et al., 2021; Xie et al., 2022), and contrastive learning (Chen et al., 2020; He et al., 2020).
Recent contrastive methods such as SimCLR (Chen et al., 2020) and MoCo (He et al., 2020) learn representations by aligning features from different augmented views, while BYOL (Grill et al., 2020) and Barlow Twins (Zbontar et al., 2021) reduce the reliance on explicit negative samples or introduce redundancy reduction.
2.2 Robust Representations in Medical Imaging
Clinical MRI segmentation tasks face challenges when transferring models to new hospitals or protocols. Public benchmarks often involve a small set of consistent sequences, limiting models to scenarios where training and testing domains match (e.g. T1w-only). Real-world deployment must handle diverse sequences and acquisition conditions.
Existing domain adaptation methods typically require multiple unlabelled images (Wang et al., 2021) or prior knowledge of the target domain (Dorent et al., 2023), which is not always feasible. SynthSeg (Billot et al., 2023) addresses this by randomising tissue contrast with synthetic data, with subsequent work showing the transferrability of the learned representations to new tasks (Liu et al., 2024a, b). The performance of these methods is expected to be constrained by the quality of the synthetic data, which may fail to capture specific anatomical details desired in downstream tasks. Similarly, Meyer et al. (2021) adjust contrast on specific regions in real images, but this approach is restricted to modest in-domain variations rather than full sequence simulation. Meanwhile, Roschewitz et al. (2024) demonstrate that generating counterfactual views can boost domain robustness for 2D chest X-ray encoders. We extend these insights to 3D MRI for sequence-invariant representations.
2.3 Quantitative MRI
Quantitative MRI (qMRI) acquires per-voxel parameter maps (e.g., $R_{1}$, $R_{2}^{*}$) that govern the signal formation in conventional scans (Weiskopf et al., 2021). These maps facilitate the simulation of numerous synthetic MRI sequences from a single qMRI acquisition (Tanenbaum et al., 2017), improving model robustness under domain shift. For example, synthesised multi-contrast data has led to enhanced results in healthy brain parcellation (Borges et al., 2021) and pathology segmentation (Chalcroft et al., 2024). Other methods rely on MR fingerprinting (Ma et al., 2013) to derive similar quantitative maps (Adams et al., 2024), further expanding opportunities for sequence-invariant learning.
3 Methods
We propose a sequence-invariant self-supervised learning (SSL) approach for robust 3D MRI representations. Figure 1 gives an overview of the method, which comprises:
  1. 1.
    A backbone encoder trained via contrastive learning on augmented volume pairs;
  2. 2.
    A decoder to reconstruct the original images, promoting semantic consistency for dense tasks;
  3. 3.
    A physics-based generative model to simulate multiple MRI sequences from quantitative MRI (qMRI) parameter maps, ensuring sequence invariance.
[…]
Figure 1: Overview of the proposed SSL approach. (1) An MPRAGE volume is augmented into two random views. We extract a feature vector $h$ via the backbone encoder, project it to $z$ for a contrastive loss $\mathcal{L}_{\text{contrastive}}$, and use a decoder to optimise a reconstruction/inpainting loss $\mathcal{L}_{\text{recon}}$. (1) We generalise this by simulating multiple scanner sequences from qMRI parameter maps, enabling sequence-invariant learning.
3.1 Self-Supervised Learning
We adopt SimCLR (Chen et al., 2020) as our core contrastive framework, though other SSL methods could also be used (Zbontar et al., 2021; Bardes et al., 2022; Oquab et al., 2024). Following Tang et al. (2022), we incorporate an additional reconstruction branch. Specifically:
  • Contrastive branch: We create two augmented 3D views of a single input volume. Each view is passed through a shared encoder, producing latent vectors $(z_{i},z_{j})$. A contrastive loss encourages $z_{i}$ and $z_{j}$ to be similar while remaining distinct from other samples in the batch. This step induces a rich feature representation that generalises well across domains.
  • Reconstruction branch: A lightweight decoder reconstructs the original volume from the latent features after removing artificially added artefacts (e.g., noise, dropout). An $L_{1}$ loss enforces pixel-level fidelity.
For spatial augmentations, each volume is randomly cropped, rotated, sheared, and flipped. We then apply MRI-specific augmentations such as non-uniform intensity fields (Van Leemput et al., 1999), Gibbs artefacts (Morelli et al., 2011), Rician noise (Gudbjartsson and Patz, 1995), and random cuboid dropout (Pathak et al., 2016). In the baseline version, we generate these augmented views from simulated MPRAGE images. In our sequence-invariant framework, we instead use parameter maps to simulate diverse MRI sequences (Sec. 3.2), enabling the encoder to learn anatomy-centric features rather than sequence-specific contrast. We train a model SeqAug that generates two views from a single simulated sequence, and a second model SeqInv that generates the two views from distinct sequence simulations, formally encouraging invariance to choice of MRI sequence.
3.2 Physics-Based Generative Model
We leverage qMRI maps (PD, $R_{1}$, $R_{2}^{*}$, MT) to synthesise multiple MRI contrasts from a single subject. Each voxel’s tissue parameters are passed through forward models approximating various standard MRI sequences (FSE, GRE, FLAIR, MPRAGE). Full signal equations are derived from known relaxation properties (Appendix A), and Rician noise is added for realism. By sampling different scanner parameters (e.g., echo time, flip angle), we obtain a range of synthetic images sharing identical anatomical structure but differing in appearance. All simulations use the NITorch library.***https://github.com/balbasty/nitorch
4 Experiments and Results
We evaluate our sequence-invariant approach on three downstream tasks: healthy brain segmentation, stroke lesion segmentation, and MRI denoising. Following standard practice, we measure segmentation performance using the Dice Similarity Coefficient (DSC) and 95th percentile Hausdorff Distance (HD95), and denoising performance using Peak Signal-to-Noise Ratio (PSNR).
4.1 Implementation Details and Data Setup
Pretraining.
We train a 3D CNN encoder (architecture detailed in Appendix B) using three different pretraining strategies:
  • Base: Trained on real MPRAGE scans only, using standard contrastive augmentations
  • SeqAug: Trained on simulated sequence images, generating two views from the same sequence
  • SeqInv: Generating two views from distinct sequence simulations, explicitly enforcing sequence invariance
All models use the NT-Xent loss (Chen et al., 2020) with temperature $\tau=0.5$ and an optional reconstruction loss with equal weighting. The pretraining dataset consists of 51 qMRI volumes (22 healthy, 29 stroke subjects), with sequence simulation performed using Bloch equations for SeqAug and SeqInv.
Downstream Tasks.
Once pretraining is complete, we freeze the encoder and optimise a U-Net decoder for:
  • Healthy Brain Segmentation: T1w, T2w, and PDw volumes from the IXI dataset (Robinson et al., 2010), segmented into background, grey matter, white matter, and CSF. We train on $96^{3}$ patches with affine and intensity augmentations, using a combined Dice + cross-entropy loss. For training, a maximum of 226 subjects are available from the GST site, with 31 reserved for validation and a further 65 for the in-domain test set. For out-of-domain testing, there are 185 and 74 subjects available in the HH and IOP sites respectively.
  • Stroke Lesion Segmentation: T1w, T2w, and FLAIR from the ARC dataset (Gibson et al., 2024). Lesions are often small, so we employ higher class weighting. We use $96^{3}$ patches and the same augmentations, optimising a combined Dice + cross-entropy loss. The T1w, T2w and FLAIR sequences are distributed in respective train/validation/test splits of (142/20/41), (159/22/47) and (59/8/18).
  • MRI Denoising: We add synthetic noise ($\sigma=0.2$) to clean IXI scans normalised to a zero mean and unit standard deviation. The network predicts the noise, which is subtracted from the input to produce the denoised image. We evaluate the result via PSNR on the same IXI splits used for healthy segmentation.
All models use $96^{3}$ patches with standard augmentations. Training details including optimization strategy, learning rate schedules, and batch sizes are provided in Appendix B.
4.2 Evaluation Metrics
Peak Signal-to-Noise Ratio (PSNR)
Assesses image quality by comparing the maximum possible signal with the noise. For an image with maximum pixel value $L$, $\text{PSNR}=20\cdot\log_{10}\bigl{(}\tfrac{L}{\sqrt{\text{MSE}}}\bigr{)}$, where $\text{MSE}=\tfrac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ measures the average error between predicted and ground truth images.
Dice Similarity Coefficient (DSC)
Measures overlap $\text{DSC}(Y,\hat{Y})=\tfrac{2|Y\cap\hat{Y}|}{|Y|+|\hat{Y}|}$ between a predicted segmentation $\hat{Y}$ and ground truth $Y$. Values range from 0 (no overlap) to 1 (perfect overlap).
95% Hausdorff Distance (HD95)
Reflects boundary accuracy by measuring the 95th percentile of all directed distances between segmentation boundaries. Smaller values indicate better delineation of anatomical edges.
4.3 Quantitative Results
4.3.1 Healthy Brain Segmentation
Table 1 compares DSC scores for T1w, T2w, and PDw images from the IXI dataset, with varying training data proportions (1%, 10%, 100% of the total available data). Our sequence-invariant (SeqInv) model consistently outperforms the baseline (Base), especially in low-data settings and out-of-domain sites (HH, IOP). Meanwhile, the sequence-augmented (SeqAug) model provides moderate gains, particularly on T2w.
Table 1: Healthy brain tissue segmentation performance using Dice Similarity Coefficient (higher is better). Values show mean ± std, with bold and underlined indicating best and second-best results. GST represents the training domain.
1% Training Data 10% Training Data 100% Training Data
Base SeqAug SeqInv Base SeqAug SeqInv Base SeqAug SeqInv
In Domain
GST [T1w] 55.1 ± 11.2 38.5 ± 12.0 56.0 ± 12.0 69.3 ± 8.2 67.2 ± 10.2 67.9 ± 9.8 89.6 ± 4.2 84.1 ± 8.5 85.5 ± 6.5
GST [T2w] 65.4 ± 6.3 56.9 ± 7.6 47.7 ± 11.2 84.2 ± 4.7 79.0 ± 4.8 68.6 ± 8.6 90.1 ± 3.4 90.5 ± 3.2 90.0 ± 3.5
GST [PDw] 38.1 ± 16.6 46.4 ± 15.9 46.6 ± 12.2 74.9 ± 9.0 70.8 ± 12.8 69.4 ± 11.2 90.1 ± 4.7 89.5 ± 5.1 90.1 ± 4.7
Out of Domain
HH [T1w] 49.4 ± 14.1 33.0 ± 14.2 57.7 ± 13.8 63.0 ± 11.7 59.3 ± 12.0 61.1 ± 13.2 81.6 ± 6.5 75.5 ± 10.9 77.4 ± 8.9
HH [T2w] 58.6 ± 7.9 53.8 ± 6.7 46.5 ± 7.0 75.0 ± 9.1 72.0 ± 8.1 65.6 ± 7.9 87.2 ± 6.4 89.7 ± 5.8 88.1 ± 6.1
HH [PDw] 33.8 ± 19.3 39.4 ± 16.2 40.3 ± 15.0 60.5 ± 15.0 61.5 ± 13.4 59.7 ± 16.4 82.7 ± 10.0 83.1 ± 10.5 85.6 ± 8.5
IOP [T1w] 50.6 ± 19.5 30.7 ± 17.9 54.4 ± 15.5 58.3 ± 15.8 60.9 ± 18.0 57.4 ± 20.1 79.1 ± 13.2 70.7 ± 16.4 74.0 ± 13.5
IOP [T2w] 58.3 ± 9.4 43.8 ± 8.3 40.6 ± 12.8 74.7 ± 6.1 71.4 ± 5.9 63.6 ± 10.0 85.1 ± 4.8 85.8 ± 5.1 86.1 ± 4.4
IOP [PDw] 31.1 ± 20.8 36.6 ± 21.3 37.6 ± 16.4 59.2 ± 13.5 55.3 ± 16.2 59.7 ± 14.1 76.4 ± 10.5 76.3 ± 11.5 77.2 ± 9.9
4.3.2 Stroke Lesion Segmentation
We next evaluate on the ARC dataset (Gibson et al., 2024) using both DSC and HD95 (see Table 2). SeqInv achieves the best overall performance on T1w, improving DSC by 0.5 points and reducing HD95 by 5.9 mm compared to the baseline. On T2w, SeqAug reduces HD95 by 22.2 mm, indicating excellent boundary accuracy while maintaining a competitive DSC. For FLAIR, SeqInv provides a further 4.7 mm decrease in HD95, offering improved boundary delineation over the baseline.
Table 2: Stroke lesion segmentation performance using 100% training data. Values show mean ± std, with bold and underlined indicating best and second-best results for each metric. DSC (higher is better) and HD95 in mm (lower is better) are shown for each model.
DSC ($\uparrow$) HD95 (mm) ($\downarrow$)
Base SeqAug SeqInv Base SeqAug SeqInv
ARC [T1w] 78.4 ± 12.8 77.3 ± 14.6 78.9 ± 12.2 33.2 ± 26.4 36.3 ± 31.0 27.3 ± 23.7
ARC [T2w] 78.7 ± 11.2 80.3 ± 9.4 79.4 ± 11.1 36.2 ± 26.5 14.0 ± 12.7 24.5 ± 24.8
ARC [FLAIR] 68.4 ± 26.7 71.0 ± 22.5 71.1 ± 23.0 67.9 ± 20.5 68.1 ± 18.4 63.2 ± 14.0
4.3.3 MRI Denoising
Lastly, we evaluate PSNR on IXI volumes corrupted with synthetic noise (Table 3). SeqInv achieves notable gains on T1w, boosting PSNR by up to 4.2 dB with only 1% training data, and these gains persist even at 100% training data, suggesting robust feature learning. Out-of-domain generalisation is also particularly strong, with SeqInv reaching 21.7 dB on HH T1w compared to 19.3 dB for the baseline. By contrast, SeqAug provides moderate gains, indicating that purely contrast-based augmentation alone cannot match the full sequence-invariant approach.
Table 3: Image denoising performance using Peak Signal-to-Noise Ratio in dB (higher is better). Values show mean ± std, with bold and underlined indicating best and second-best results. GST represents the training domain.
1% Training Data 10% Training Data 100% Training Data
Base SeqAug SeqInv Base SeqAug SeqInv Base SeqAug SeqInv
In Domain
GST [T1w] 14.9 ± 0.2 16.2 ± 0.3 19.1 ± 0.4 19.0 ± 0.4 19.7 ± 0.8 20.3 ± 1.0 19.1 ± 0.3 20.6 ± 0.4 21.3 ± 0.4
GST [T2w] 17.2 ± 0.2 17.7 ± 0.2 17.3 ± 0.3 18.3 ± 0.3 18.5 ± 0.6 19.8 ± 0.3 18.3 ± 0.3 19.4 ± 0.3 20.0 ± 0.4
GST [PDw] 17.0 ± 0.2 18.3 ± 0.3 18.7 ± 0.3 18.6 ± 0.3 19.3 ± 0.5 20.0 ± 0.5 18.7 ± 0.3 19.9 ± 0.3 20.6 ± 0.3
Out of Domain
HH [T1w] 15.1 ± 0.1 16.5 ± 0.2 19.4 ± 0.2 19.1 ± 0.3 20.0 ± 0.7 20.1 ± 1.1 19.3 ± 0.2 21.0 ± 0.3 21.7 ± 0.3
HH [T2w] 16.5 ± 0.4 16.9 ± 0.5 16.4 ± 0.5 17.5 ± 0.4 15.6 ± 1.4 18.8 ± 0.6 17.5 ± 0.4 18.5 ± 0.6 18.9 ± 0.6
HH [PDw] 16.5 ± 0.4 17.8 ± 0.5 17.8 ± 0.6 18.0 ± 0.5 19.2 ± 0.6 18.9 ± 1.0 18.2 ± 0.5 19.3 ± 0.6 19.9 ± 0.7
IOP [T1w] 14.7 ± 0.2 16.7 ± 0.2 18.9 ± 0.2 18.4 ± 0.2 19.9 ± 0.3 18.5 ± 0.9 18.8 ± 0.1 20.3 ± 0.2 21.0 ± 0.2
IOP [T2w] 17.1 ± 0.1 17.6 ± 0.1 17.0 ± 0.1 17.9 ± 0.2 18.8 ± 0.3 19.6 ± 0.2 18.0 ± 0.1 19.2 ± 0.1 19.7 ± 0.2
IOP [PDw] 16.9 ± 0.1 18.3 ± 0.1 18.8 ± 0.1 18.5 ± 0.1 19.8 ± 0.2 20.0 ± 0.3 18.6 ± 0.1 19.8 ± 0.1 20.5 ± 0.2
5 Discussion
Our results show that sequence-invariant self-supervised learning substantially improves model robustness and generalisation across diverse MRI sequences and acquisition sites. In particular, it enables effective feature learning even with minimal labelled data, suggesting that the method captures fundamental anatomical cues independent of sequence-specific contrast.
5.1 Key Findings
We highlight three key aspects of our method’s performance. First, even when trained on as little as 1% of the data, it achieves up to +4.2 dB PSNR in denoising and +8.3 DSC points in segmentation, underscoring its robust representation capabilities. Second, the model generalises well across T1w, T2w, and PDw, showing particularly strong results on T1w while leaving some gaps on the other sequences. Finally, it excels at out-of-domain adaptation, often surpassing baseline models more in unseen sites than in the original training domain, illustrating its effectiveness for cross-site generalisation.
5.2 Limitations
Our approach faces several limitations. First, training on full-resolution 3D volumes is computationally demanding, limiting batch size and reducing the number of negative pairs used during contrastive learning. Second, we rely on a CNN backbone, which may not capture long-range dependencies as effectively as vision transformers or other recent architectures. Third, while cross-sequence invariance bolsters model robustness, certain sequence-specific gaps - particularly on T2w images - highlight the need for further improvements. Finally, qMRI data are relatively scarce, leading us to pre-train on only 51 subjects; scaling to a larger qMRI dataset or inferring parameter maps from a resource like the UK Biobank (Sudlow et al., 2015) could significantly boost downstream performance.
5.3 Future Directions
We plan to explore ViT-based encoders, alternative SSL methods (VICReg, DINO), and larger clinical datasets to further improve domain invariance and push towards routine clinical adoption. We also expect multi-view contrastive learning and decoder pretraining (Asiedu et al., 2022) to be valuable directions of future work.
5.4 Conclusion
Sequence-invariant self-supervised learning offers a promising route towards more robust, generalisable medical image analysis. By using physics-informed contrast simulation and contrastive training, we can exploit the shared anatomy across varied MRI sequences and sites. Although challenges remain - especially around computational cost and data availability - our results illustrate the potential for significant gains in low-data scenarios and out-of-domain adaptation. We believe this framework provides a stepping stone toward truly cross-domain, clinically deployable deep learning models in medical imaging.
[…]
LC is supported by the EPSRC-funded UCL Centre for Doctoral Training in Intelligent, Integrated Imaging in Healthcare (i4health) (EP/S021930/1), and the Wellcome Trust (203147/Z/16/Z and 205103/Z/16/Z). This research was supported by NVIDIA and used NVIDIA RTX A6000 48GB.
Appendix A Physics-based Signal Equations
This appendix details the forward models used to simulate MRI signal intensities from quantitative MRI (qMRI) parameter maps. For each voxel, we assume known proton density (PD), longitudinal relaxation rate ($R_{1}$), and transverse relaxation rate ($R_{2}$ or $R_{2}^{*}$), and optionally a magnetic transfer (MT) parameter. We consider several common MRI sequences, each governed by its own signal equation:
Fast Spin-Echo (FSE).
$S_{\text{FSE}}=B_{1}\cdot\text{PD}\cdot\bigl{(}1-e^{-R_{1}\cdot T_{R}}\bigr{)}\cdot e^{-R_{2}\cdot T_{E}}.$ (1)
Gradient-Echo (GRE).
$S_{\text{GRE}}=B_{1}\cdot\text{PD}\cdot\sin(\alpha)\cdot(1-\text{MT})\cdot\tfrac{1-e^{-R_{1}\cdot T_{R}}}{1-\cos(\alpha)(1-\text{MT})\,e^{-R_{1}\cdot T_{R}}}\cdot e^{-R_{2}^{*}\cdot T_{E}}.$ (2)
Fluid-Attenuated Inversion Recovery (FLAIR).
$S_{\text{FLAIR}}=B_{1}\cdot\text{PD}\cdot e^{-R_{2}\cdot T_{E}}\cdot\bigl{(}1-2\,e^{-R_{1}\cdot T_{I}}+e^{-R_{1}\cdot T_{R}}\bigr{)}.$ (3)
Magnetisation-Prepared Rapid Gradient Echo (MPRAGE).
$S_{\text{MPRAGE}}$ $=\Bigl{|}\sin(\alpha)\,\Bigl{[}\tfrac{1-e^{-R_{1}\cdot T_{R}}}{1-\cos(\alpha)\,e^{-R_{1}\cdot T_{R}}}\,\bigl{(}1-(\cos(\alpha)\,e^{-T_{X}\cdot R_{1}})^{n}\bigr{)}\Bigr{]}\,e^{-T_{D}\cdot R_{1}}$ (4)
$\quad+\,\bigl{(}1-e^{-T_{D}\cdot R_{1}}\bigr{)}\Bigr{|}.$
Here, $B_{1}$ represents the receive field strength, $\alpha$ (in radians) is the flip angle, $T_{R}$ the repetition time, $T_{E}$ the echo time, $T_{I}$ the inversion time, $T_{X}$ the excitation repetition time, $T_{D}$ a delay time, and $n$ the number of excitation pulses (all times in seconds unless noted otherwise).
Noise Simulation
To introduce realistic MRI noise, we add Rician noise Gudbjartsson and Patz (1995) to the simulated signals. Specifically,
$S_{\textrm{noisy}}=\sqrt{(S_{\textrm{MRI}}+\epsilon_{\textrm{r}})^{2}+\epsilon_{\textrm{i}}^{2}}\text{ where }\epsilon_{\textrm{r}}\sim\mathcal{N}(0,\sigma^{2})\text{ and }\epsilon_{\textrm{i}}\sim\mathcal{N}(0,\sigma^{2})$ (5)
where $S_{\text{MRI}}$ is the noise-free signal, and each $\mathcal{N}(0,\sigma^{2})$ term is a zero-mean Gaussian simulating real and imaginary noise components.
All simulations were implemented using the NiTorch libraryhttps://github.com/balbasty/nitorch.
Acquisition Parameters and Data Augmentation
We randomly sampled MRI sequences and parameters from realistic distributions to produce diverse training data, as summarised below. Each sequence was chosen with equal probability to ensure broad coverage of acquisition protocols.
FLAIR:
$T_{E}\sim 10^{\mathcal{N}(\log_{10}(0.02),\log_{10}(0.1))}\quad\text{s},\quad T_{R}\sim 10^{\mathcal{U}(\log_{10}(0.001),\log_{10}(5))}\quad\text{s},$
$T_{I}\sim 10^{\mathcal{U}(\log_{10}(0.001),\log_{10}(3))}\quad\text{s}.$
FSE:
$T_{E}\sim 10^{\mathcal{U}(\log_{10}(0.001),\log_{10}(3))}\quad\text{s},\quad T_{R}\sim 10^{\mathcal{U}(\log_{10}(0.001),\log_{10}(3))}\quad\text{s}.$
MPRAGE:
$T_{R}\sim\mathcal{N}(23,2.3)\quad\text{s},\quad T_{I}\sim\mathcal{U}(0.6,0.9)\quad\text{s},\quad T_{X}\sim\mathcal{U}(0.004,0.008)\quad\text{s},$
$T_{E}\sim\mathcal{U}(0.002,0.004)\quad\text{s},\quad\alpha\sim\mathcal{U}(5^{\circ},12^{\circ}).$
GRE:
$T_{E}\sim 10^{\mathcal{U}(\log_{10}(0.002),\log_{10}(0.08))}\quad\text{s},\quad T_{R}\sim 10^{\mathcal{U}(\log_{10}(0.005),\log_{10}(5))}\quad\text{s},\quad\alpha\sim\mathcal{U}(5^{\circ},50^{\circ}).$
All sampled parameters were constrained to physically plausible ranges (e.g. taking absolute values as needed). This broad variation in sequences and parameter values helps train the model to handle diverse imaging protocols.
Appendix B Model Architectures
Pre-training Architecture
The pre-training setup consists of three components:
  • CNN Encoder:
    • Input: 3D volume with 1 channel
    • 5 downsampling blocks with features: (64, 128, 256, 512, 768)
    • Each block: Two 3D convolutions with kernel size 3
    • Instance normalization and GELU activation after each convolution
    • Dropout rate: 0.2
  • Projector: Two-layer MLP
    • Input dimension: 768
    • Hidden layer (768 → 512) with GELU activation
    • Output projects to contrastive learning space (512 → 128)
  • Reconstructor: Decoder network
    • 4 transposed convolution blocks: (768 → 384 → 192 → 96 → 48)
    • Each upsampling: 2×2×2 stride
    • Final 1×1×1 convolution to output channels
Downstream Task Architectures
For the denoising task, we use a U-Net architecture that incorporates the pre-trained encoder:
  • CNN U-Net:
    • Input: 3D volume with 1 channel
    • Encoder: Pre-trained CNN encoder (frozen)
    • Feature dimensions: (768, 512, 256, 128, 64, 32)
    • Instance normalization throughout
    • GELU activation functions
    • Dropout rate: 0.2
    • Upsampling: Transposed convolutions
    • Output: 1 channel (predicted noise)
B.1 Training Details
The models were trained with the following specifications:
  • Optimizer: AdamW with gradient clipping at 12.0
  • Learning rate schedule: $(1-\frac{epoch}{max\_epochs})^{0.9}$
  • Loss functions:
    • Pre-training: NT-Xent loss + L1 reconstruction loss
    • Denoising: Mean Squared Error (MSE)
    • Segmentation: Dice + Cross-Entropy
  • Patch size: 96×96×96
  • Mixed precision training
  • Batch size:
    • Pre-training: 8
    • Downstream tasks: 2
During downstream task training, the pre-trained encoder weights were frozen while the decoder weights were trained from scratch, as evidenced by the weight loading and gradient freezing in the training code.

### v3

@@
-
+Unified 3D MRI Representations via Sequence-Invariant Contrastive Learning
 Abstract
@@
 By simulating multiple MRI contrasts from a single 3D qMRI scan and enforcing consistent representations across these contrasts, we learn anatomy-centric rather than sequence-specific features.
-This yields a robust 3D encoder that performs strongly across varied tasks and protocols.
+The result is a single 3D encoder that excels across tasks and protocols.
 Experiments on healthy brain segmentation (IXI), stroke lesion segmentation (ARC), and MRI denoising show significant gains over baseline SSL approaches, especially in low-data settings (up to +8.3% Dice, +4.2 dB PSNR).
-Our model also generalises effectively to unseen sites, demonstrating potential for more scalable and clinically reliable volumetric analysis.
-All code and trained models are publicly available.
-editors: Under Review for MIDL 2025
+It also generalises to unseen sites, supporting scalable clinical use.
+Code and trained models are publicly available.
 1 Introduction
-Deep learning has advanced medical image analysis across a range of tasks, from image registration [Balakrishnan et al.
-(2019)] to semantic segmentation [Chalcroft et al.
-(2023)] and visual question-answering [Narayanan et al.
-(2024)].
-However, unique challenges arise when working with 3D MRI data, including increased computational demands and the difficulty of applying 2D pre-trained models to volumetric contexts [Ma et al.
-(2024)].
-Although large-scale 3D datasets [Li et al.
-(2024)] and models [Wu et al.
-(2024)] have recently emerged, fine-tuning them for specific clinical tasks remains non-trivial due to inevitable domain shifts [Yang et al.
-(2024)].
-Self-supervised learning (SSL) offers a promising means of learning robust representations without the need for large labelled datasets [Chaitanya et al.
-(2020), Huang et al.
-(2023)].
+Deep learning now underpins medical image registration [3] and segmentation [8].
+However, unique challenges arise when working with 3D MRI data, including increased computational demands and the difficulty of applying 2D pre-trained models to volumetric contexts [17].
+Although large-scale 3D datasets and models [27] have recently emerged, fine-tuning them for specific clinical tasks remains non-trivial due to inevitable domain shifts [28].
+Self-supervised learning (SSL) offers a promising means of learning robust representations without the need for large labelled datasets.
 Yet, existing SSL methods often treat each MRI sequence as a separate domain, neglecting the shared anatomical information across contrast variations.
 In contrast, we leverage the observation that different MRI sequences, despite their unique contrast properties, encode the same underlying anatomy.
-By incorporating physics-based knowledge of MRI signal generation into an SSL framework, we can learn representations that are robust to sequence changes while preserving anatomical fidelity.
-In this paper, we make the following contributions:
-  • We propose a sequence-invariant self-supervised learning framework based on physics-driven contrast simulation, enabling robust 3D MRI representations across multiple sequences.
-  • We show significant performance gains in low-data regimes, with up to 8.3% Dice improvement in segmentation and a 4.2 dB PSNR increase for denoising when using just 1% of the training data.
-  • We demonstrate enhanced cross-sequence generalisation and site adaptation from a single pre-trained encoder, particularly in out-of-domain scenarios.
-  • We provide comprehensive evaluations on three diverse tasks - healthy brain segmentation, stroke lesion segmentation, and image denoising - highlighting the clinical utility of our approach.
+Embedding MRI physics in SSL is expected to yield representations that ignore sequence contrast yet keep anatomy.
+We (i) introduce a physics-driven, sequence-invariant SSL framework, (ii) boost Dice by up to 8.3% and PSNR by 4.2 dB with only 1% labels, and (iii) show strong cross-site generalisation.
+We provide comprehensive evaluations of the prosed method on three diverse tasks - healthy brain segmentation, stroke lesion segmentation, and image denoising - highlighting the clinical utility of our approach.
 Our method addresses key problems in medical imaging by enabling robust feature learning across different sites and sequences, even with limited annotated data.
@@
 Self-supervised learning (SSL) can leverage unlabelled data by creating proxy tasks that encourage useful invariances in learned representations.
-Techniques include predictive coding (van den Oord et al., 2019; Hénaff et al., 2020), masked image modelling (He et al., 2021; Xie et al., 2022), and contrastive learning (Chen et al., 2020; He et al., 2020).
-Recent contrastive methods such as SimCLR (Chen et al., 2020) and MoCo (He et al., 2020) learn representations by aligning features from different augmented views, while BYOL (Grill et al., 2020) and Barlow Twins (Zbontar et al., 2021) reduce the reliance on explicit negative samples or introduce redundancy reduction.
+Techniques include predictive coding, masked image modelling, and contrastive learning.
+Recent contrastive methods such as SimCLR [9] and MoCo [15] learn representations by aligning features from different augmented views, while BYOL [12] and Barlow Twins [29] reduce the reliance on explicit negative samples or introduce redundancy reduction.
+SSL is now routine in medical imaging for using unlabelled data to boost downstream tasks.
+Adopted methods include contrastive learning [24], masked image modelling [25] and reconstruction-based proxy tasks [19, 30].
 2.2 Robust Representations in Medical Imaging
@@
 Real-world deployment must handle diverse sequences and acquisition conditions.
-Existing domain adaptation methods typically require multiple unlabelled images (Wang et al., 2021) or prior knowledge of the target domain (Dorent et al., 2023), which is not always feasible.
-SynthSeg (Billot et al., 2023) addresses this by randomising tissue contrast with synthetic data, with subsequent work showing the transferrability of the learned representations to new tasks (Liu et al., 2024a, b).
-The performance of these methods is expected to be constrained by the quality of the synthetic data, which may fail to capture specific anatomical details desired in downstream tasks.
-Similarly, Meyer et al.
-(2021) adjust contrast on specific regions in real images, but this approach is restricted to modest in-domain variations rather than full sequence simulation.
-Meanwhile, Roschewitz et al.
-(2024) demonstrate that generating counterfactual views can boost domain robustness for 2D chest X-ray encoders.
+Existing domain adaptation methods typically require multiple unlabelled images or prior knowledge of the target domain [10], which is not always feasible.
+SynthSeg [4] addresses this by randomising tissue contrast with synthetic data, with subsequent work showing the transferrability of the learned representations to new tasks.
+Their success hinges on synthetic data quality, which may miss fine anatomy.
+Similarly, [18] adjust contrast on specific regions in real images, but this approach is restricted to modest in-domain variations rather than full sequence simulation.
+Meanwhile, [22] demonstrate that generating counterfactual views can boost domain robustness for 2D chest X-ray encoders.
 We extend these insights to 3D MRI for sequence-invariant representations.
 2.3 Quantitative MRI
-Quantitative MRI (qMRI) acquires per-voxel parameter maps (e.g., $R_{1}$, $R_{2}^{*}$) that govern the signal formation in conventional scans (Weiskopf et al., 2021).
-These maps facilitate the simulation of numerous synthetic MRI sequences from a single qMRI acquisition (Tanenbaum et al., 2017), improving model robustness under domain shift.
-For example, synthesised multi-contrast data has led to enhanced results in healthy brain parcellation (Borges et al., 2021) and pathology segmentation (Chalcroft et al., 2024).
-Other methods rely on MR fingerprinting (Ma et al., 2013) to derive similar quantitative maps (Adams et al., 2024), further expanding opportunities for sequence-invariant learning.
+Quantitative MRI (qMRI) acquires per-voxel parameter maps (e.g., $R_{1}$, $R_{2}^{*}$) that govern the signal formation in conventional scans [26].
+These maps facilitate the simulation of numerous synthetic MRI sequences from a single qMRI acquisition [23], improving model robustness under domain shift.
+For example, synthesised multi-contrast data has led to enhanced results in healthy brain parcellation [5], improved visualisation and segmentation of subcortical structures through synthetic multi-inversion-time contrasts [14], and better pathology segmentation [7].
+Other methods rely on MR fingerprinting [16] to derive similar quantitative maps [1], further expanding opportunities for sequence-invariant learning.
 3 Methods
-We propose a sequence-invariant self-supervised learning (SSL) approach for robust 3D MRI representations. Figure 1 gives an overview of the method, which comprises:
-  1. 1.
-    A backbone encoder trained via contrastive learning on augmented volume pairs;
-  2. 2.
-    A decoder to reconstruct the original images, promoting semantic consistency for dense tasks;
-  3. 3.
-    A physics-based generative model to simulate multiple MRI sequences from quantitative MRI (qMRI) parameter maps, ensuring sequence invariance.
-[…]
-[…]
+We propose sequence-invariant SSL for robust 3D MRI representations.
+Figure 1 shows (i) a contrastive encoder, (ii) a reconstruction decoder, and (iii) a physics engine that simulates multiple sequences from qMRI.
 Figure 1: Overview of the proposed SSL approach.
-(1) An MPRAGE volume is augmented into two random views.
+(1(a)) Baseline: An MPRAGE volume is augmented into two random views.
 We extract a feature vector $h$ via the backbone encoder, project it to $z$ for a contrastive loss $\mathcal{L}_{\text{contrastive}}$, and use a decoder to optimise a reconstruction/inpainting loss $\mathcal{L}_{\text{recon}}$.
-(1) We generalise this by simulating multiple scanner sequences from qMRI parameter maps, enabling sequence-invariant learning.
+(1(b)) SeqAug/SeqInv: We generalise this by simulating multiple scanner sequences from qMRI parameter maps, enabling sequence-invariant learning.
+In SeqAug both $\tau$, $\tau^{\prime}$ would produce augmented views of the same sequence, while in SeqInv both views will contain a different sequence.
 3.1 Self-Supervised Learning
-We adopt SimCLR (Chen et al., 2020) as our core contrastive framework, though other SSL methods could also be used (Zbontar et al., 2021; Bardes et al., 2022; Oquab et al., 2024).
-Following Tang et al.
-(2022), we incorporate an additional reconstruction branch.
-Specifically:
+We adopt SimCLR [9] as our core contrastive framework, though other SSL methods could also be used. Following [24], we incorporate an additional reconstruction branch. Specifically:
   • Contrastive branch: We create two augmented 3D views of a single input volume.
@@
 An $L_{1}$ loss enforces pixel-level fidelity.
-For spatial augmentations, each volume is randomly cropped, rotated, sheared, and flipped.
-We then apply MRI-specific augmentations such as non-uniform intensity fields (Van Leemput et al., 1999), Gibbs artefacts (Morelli et al., 2011), Rician noise (Gudbjartsson and Patz, 1995), and random cuboid dropout (Pathak et al., 2016).
+Spatial augmentations include random crops, rotations, shears and flips.
+We then apply MRI-specific augmentations such as non-uniform intensity fields, Gibbs artefacts, Rician noise and random cuboid dropout [20].
 In the baseline version, we generate these augmented views from simulated MPRAGE images.
+Magnetisation-Prepared RApid Gradient Echo (MPRAGE) is a common T1-weighted structural MRI sequence, particularly common in research studies.
 In our sequence-invariant framework, we instead use parameter maps to simulate diverse MRI sequences (Sec.
@@
 We train a model SeqAug that generates two views from a single simulated sequence, and a second model SeqInv that generates the two views from distinct sequence simulations, formally encouraging invariance to choice of MRI sequence.
-3.2 Physics-Based Generative Model
+Our baseline model (Base) was pretrained exclusively using synthetic MPRAGE images generated from qMRI parameter maps, ensuring a fair comparison to our proposed methods.
+3.2 Physics-Based Data Synthesis
 We leverage qMRI maps (PD, $R_{1}$, $R_{2}^{*}$, MT) to synthesise multiple MRI contrasts from a single subject.
 Each voxel’s tissue parameters are passed through forward models approximating various standard MRI sequences (FSE, GRE, FLAIR, MPRAGE).
-Full signal equations are derived from known relaxation properties (Appendix A), and Rician noise is added for realism.
+Full signal equations are derived from known relaxation properties (Appendix 0.A), and Rician noise is added for realism.
 By sampling different scanner parameters (e.g., echo time, flip angle), we obtain a range of synthetic images sharing identical anatomical structure but differing in appearance.
-All simulations use the NITorch library.***https://github.com/balbasty/nitorch
+All simulations use the NITorch library [^1] https://github.com/balbasty/nitorch.
 4 Experiments and Results
@@
 Pretraining.
-We train a 3D CNN encoder (architecture detailed in Appendix B) using three different pretraining strategies:
-  • Base: Trained on real MPRAGE scans only, using standard contrastive augmentations
-  • SeqAug: Trained on simulated sequence images, generating two views from the same sequence
-  • SeqInv: Generating two views from distinct sequence simulations, explicitly enforcing sequence invariance
-All models use the NT-Xent loss (Chen et al., 2020) with temperature $\tau=0.5$ and an optional reconstruction loss with equal weighting.
+We pre-train three encoders: Base (real MPRAGE only), SeqAug (two views of one simulated sequence) and SeqInv (views from two simulated sequences). Architecture details are given in Appendix 0.B.
+All models use NT-Xent [9] (temperature 0.5) plus an equally-weighted $L_{1}$ reconstruction term.
 The pretraining dataset consists of 51 qMRI volumes (22 healthy, 29 stroke subjects), with sequence simulation performed using Bloch equations for SeqAug and SeqInv.
@@
 Once pretraining is complete, we freeze the encoder and optimise a U-Net decoder for:
-  • Healthy Brain Segmentation: T1w, T2w, and PDw volumes from the IXI dataset (Robinson et al., 2010), segmented into background, grey matter, white matter, and CSF.
+  • Healthy Brain Segmentation: T1w, T2w, and PDw volumes from the IXI dataset [21], segmented into background, grey matter, white matter, and CSF.
 We train on $96^{3}$ patches with affine and intensity augmentations, using a combined Dice + cross-entropy loss.
@@
 For out-of-domain testing, there are 185 and 74 subjects available in the HH and IOP sites respectively.
-  • Stroke Lesion Segmentation: T1w, T2w, and FLAIR from the ARC dataset (Gibson et al., 2024).
+  • Stroke Lesion Segmentation: T1w, T2w, and FLAIR from the ARC dataset [11].
 Lesions are often small, so we employ higher class weighting.
@@
 We evaluate the result via PSNR on the same IXI splits used for healthy segmentation.
-All models use $96^{3}$ patches with standard augmentations. Training details including optimization strategy, learning rate schedules, and batch sizes are provided in Appendix B.
+All models use $96^{3}$ patches with standard augmentations.
+Training details including optimization strategy, learning rate schedules, and batch sizes are provided in Appendix 0.B.
+A new decoder is trained for each task/model.
 4.2 Evaluation Metrics
@@
 For an image with maximum pixel value $L$, $\text{PSNR}=20\cdot\log_{10}\bigl{(}\tfrac{L}{\sqrt{\text{MSE}}}\bigr{)}$, where $\text{MSE}=\tfrac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ measures the average error between predicted and ground truth images.
+While PSNR clearly quantifies denoising improvements, perceptual metrics such as SSIM or LPIPS might provide better insight into human-perceived image quality and will be explored in future analyses.
 Dice Similarity Coefficient (DSC)
@@
 Meanwhile, the sequence-augmented (SeqAug) model provides moderate gains, particularly on T2w.
+In the 1% training data regime, Base performs particularly well on the T2w data; looking at the individual tissue class metrics in Table 1, it appears that this is most notable in the White Matter tissue class.
+We posit that this may be due to the SeqAug/SeqInv models’ aversions to learning sequence-specific features preventing them from easily leveraging domain-specific cues such as White Matter in T2-weighted MRI being much darker than any surrounding tissue.
 Table 1: Healthy brain tissue segmentation performance using Dice Similarity Coefficient (higher is better).
-Values show mean ± std, with bold and underlined indicating best and second-best results.
+Values show mean ± standard error, with bold and underlined indicating best and second-best results.
 GST represents the training domain.
@@
 In Domain
-GST [T1w] 55.1 ± 11.2 38.5 ± 12.0 56.0 ± 12.0 69.3 ± 8.2 67.2 ± 10.2 67.9 ± 9.8 89.6 ± 4.2 84.1 ± 8.5 85.5 ± 6.5
-GST [T2w] 65.4 ± 6.3 56.9 ± 7.6 47.7 ± 11.2 84.2 ± 4.7 79.0 ± 4.8 68.6 ± 8.6 90.1 ± 3.4 90.5 ± 3.2 90.0 ± 3.5
-GST [PDw] 38.1 ± 16.6 46.4 ± 15.9 46.6 ± 12.2 74.9 ± 9.0 70.8 ± 12.8 69.4 ± 11.2 90.1 ± 4.7 89.5 ± 5.1 90.1 ± 4.7
+GST [T1w] 55.1 ± 0.8 38.5 ± 0.9 56.0 ± 0.9 69.3 ± 0.6 67.2 ± 0.7 67.9 ± 0.7 89.6 ± 0.3 84.1 ± 0.6 85.5 ± 0.5
+GST [T2w] 65.4 ± 0.4 56.9 ± 0.5 47.7 ± 0.8 84.2 ± 0.3 79.0 ± 0.3 68.6 ± 0.6 90.1 ± 0.2 90.5 ± 0.2 90.0 ± 0.2
+GST [PDw] 38.1 ± 1.2 46.4 ± 1.1 46.6 ± 0.9 74.9 ± 0.6 70.8 ± 0.9 69.4 ± 0.8 90.1 ± 0.3 89.5 ± 0.4 90.1 ± 0.3
 Out of Domain
-HH [T1w] 49.4 ± 14.1 33.0 ± 14.2 57.7 ± 13.8 63.0 ± 11.7 59.3 ± 12.0 61.1 ± 13.2 81.6 ± 6.5 75.5 ± 10.9 77.4 ± 8.9
-HH [T2w] 58.6 ± 7.9 53.8 ± 6.7 46.5 ± 7.0 75.0 ± 9.1 72.0 ± 8.1 65.6 ± 7.9 87.2 ± 6.4 89.7 ± 5.8 88.1 ± 6.1
-HH [PDw] 33.8 ± 19.3 39.4 ± 16.2 40.3 ± 15.0 60.5 ± 15.0 61.5 ± 13.4 59.7 ± 16.4 82.7 ± 10.0 83.1 ± 10.5 85.6 ± 8.5
-IOP [T1w] 50.6 ± 19.5 30.7 ± 17.9 54.4 ± 15.5 58.3 ± 15.8 60.9 ± 18.0 57.4 ± 20.1 79.1 ± 13.2 70.7 ± 16.4 74.0 ± 13.5
-IOP [T2w] 58.3 ± 9.4 43.8 ± 8.3 40.6 ± 12.8 74.7 ± 6.1 71.4 ± 5.9 63.6 ± 10.0 85.1 ± 4.8 85.8 ± 5.1 86.1 ± 4.4
-IOP [PDw] 31.1 ± 20.8 36.6 ± 21.3 37.6 ± 16.4 59.2 ± 13.5 55.3 ± 16.2 59.7 ± 14.1 76.4 ± 10.5 76.3 ± 11.5 77.2 ± 9.9
+HH [T1w] 49.4 ± 0.6 33.0 ± 0.6 57.7 ± 0.6 63.0 ± 0.5 59.3 ± 0.5 61.1 ± 0.6 81.6 ± 0.3 75.5 ± 0.5 77.4 ± 0.4
+HH [T2w] 58.6 ± 0.3 53.8 ± 0.3 46.5 ± 0.3 75.0 ± 0.4 72.0 ± 0.3 65.6 ± 0.3 87.2 ± 0.3 89.7 ± 0.2 88.1 ± 0.3
+HH [PDw] 33.8 ± 0.8 39.4 ± 0.7 40.3 ± 0.6 60.5 ± 0.6 61.5 ± 0.6 59.7 ± 0.7 82.7 ± 0.4 83.1 ± 0.4 85.6 ± 0.4
+IOP [T1w] 50.6 ± 1.3 30.7 ± 1.2 54.4 ± 1.0 58.3 ± 1.1 60.9 ± 1.2 57.4 ± 1.3 79.1 ± 0.9 70.7 ± 1.1 74.0 ± 0.9
+IOP [T2w] 58.3 ± 0.6 43.8 ± 0.6 40.6 ± 0.9 74.7 ± 0.4 71.4 ± 0.4 63.6 ± 0.7 85.1 ± 0.3 85.8 ± 0.3 86.1 ± 0.3
+IOP [PDw] 31.1 ± 1.4 36.6 ± 1.4 37.6 ± 1.1 59.2 ± 0.9 55.3 ± 1.1 59.7 ± 0.9 76.4 ± 0.7 76.3 ± 0.8 77.2 ± 0.7
 4.3.2 Stroke Lesion Segmentation
-We next evaluate on the ARC dataset (Gibson et al., 2024) using both DSC and HD95 (see Table 2).
+We next evaluate on the ARC dataset [11] using both DSC and HD95 (see Table 2).
 SeqInv achieves the best overall performance on T1w, improving DSC by 0.5 points and reducing HD95 by 5.9 mm compared to the baseline.
@@
 Table 2: Stroke lesion segmentation performance using 100% training data.
-Values show mean ± std, with bold and underlined indicating best and second-best results for each metric.
+Values show mean ± standard error, with bold and underlined indicating best and second-best results for each metric.
 DSC (higher is better) and HD95 in mm (lower is better) are shown for each model.
-DSC ($\uparrow$) HD95 (mm) ($\downarrow$)
+DSC HD95 (mm)
 Base SeqAug SeqInv Base SeqAug SeqInv
-ARC [T1w] 78.4 ± 12.8 77.3 ± 14.6 78.9 ± 12.2 33.2 ± 26.4 36.3 ± 31.0 27.3 ± 23.7
-ARC [T2w] 78.7 ± 11.2 80.3 ± 9.4 79.4 ± 11.1 36.2 ± 26.5 14.0 ± 12.7 24.5 ± 24.8
-ARC [FLAIR] 68.4 ± 26.7 71.0 ± 22.5 71.1 ± 23.0 67.9 ± 20.5 68.1 ± 18.4 63.2 ± 14.0
+ARC [T1w] 78.4 ± 2.0 77.3 ± 2.3 78.9 ± 1.9 33.2 ± 4.1 36.3 ± 4.8 27.3 ± 3.7
+ARC [T2w] 78.7 ± 1.6 80.3 ± 1.4 79.4 ± 1.6 36.2 ± 3.9 14.0 ± 1.9 24.5 ± 3.6
+ARC [FLAIR] 68.4 ± 6.3 71.0 ± 5.3 71.1 ± 5.4 67.9 ± 4.8 68.1 ± 4.3 63.2 ± 3.3
 4.3.3 MRI Denoising
@@
 By contrast, SeqAug provides moderate gains, indicating that purely contrast-based augmentation alone cannot match the full sequence-invariant approach.
+It is notable that the SeqInv model’s benefit is much more apparent in this denoising task compared to the previous segmentation tasks.
+This could be explained by the similarity of image restoration tasks to the objective of contrastive learning to learn invariance to view augmentations.
+The heavier constraint on invariance due to view-dependent sequences may be better suited to image restoration tasks than discriminative tasks like segmentation and classification.
 Table 3: Image denoising performance using Peak Signal-to-Noise Ratio in dB (higher is better).
-Values show mean ± std, with bold and underlined indicating best and second-best results.
+Values show mean ± standard error, with bold and underlined indicating best and second-best results.
 GST represents the training domain.
@@
 In Domain
-GST [T1w] 14.9 ± 0.2 16.2 ± 0.3 19.1 ± 0.4 19.0 ± 0.4 19.7 ± 0.8 20.3 ± 1.0 19.1 ± 0.3 20.6 ± 0.4 21.3 ± 0.4
-GST [T2w] 17.2 ± 0.2 17.7 ± 0.2 17.3 ± 0.3 18.3 ± 0.3 18.5 ± 0.6 19.8 ± 0.3 18.3 ± 0.3 19.4 ± 0.3 20.0 ± 0.4
-GST [PDw] 17.0 ± 0.2 18.3 ± 0.3 18.7 ± 0.3 18.6 ± 0.3 19.3 ± 0.5 20.0 ± 0.5 18.7 ± 0.3 19.9 ± 0.3 20.6 ± 0.3
+GST [T1w] 14.9 ± 0.0 16.2 ± 0.0 19.1 ± 0.0 19.0 ± 0.1 19.7 ± 0.1 20.3 ± 0.1 19.1 ± 0.0 20.6 ± 0.0 21.3 ± 0.1
+GST [T2w] 17.2 ± 0.0 17.7 ± 0.0 17.3 ± 0.0 18.3 ± 0.0 18.5 ± 0.1 19.8 ± 0.0 18.3 ± 0.0 19.4 ± 0.0 20.0 ± 0.0
+GST [PDw] 17.0 ± 0.0 18.3 ± 0.0 18.7 ± 0.0 18.6 ± 0.0 19.3 ± 0.1 20.0 ± 0.1 18.7 ± 0.0 19.9 ± 0.0 20.6 ± 0.0
 Out of Domain
-HH [T1w] 15.1 ± 0.1 16.5 ± 0.2 19.4 ± 0.2 19.1 ± 0.3 20.0 ± 0.7 20.1 ± 1.1 19.3 ± 0.2 21.0 ± 0.3 21.7 ± 0.3
-HH [T2w] 16.5 ± 0.4 16.9 ± 0.5 16.4 ± 0.5 17.5 ± 0.4 15.6 ± 1.4 18.8 ± 0.6 17.5 ± 0.4 18.5 ± 0.6 18.9 ± 0.6
-HH [PDw] 16.5 ± 0.4 17.8 ± 0.5 17.8 ± 0.6 18.0 ± 0.5 19.2 ± 0.6 18.9 ± 1.0 18.2 ± 0.5 19.3 ± 0.6 19.9 ± 0.7
-IOP [T1w] 14.7 ± 0.2 16.7 ± 0.2 18.9 ± 0.2 18.4 ± 0.2 19.9 ± 0.3 18.5 ± 0.9 18.8 ± 0.1 20.3 ± 0.2 21.0 ± 0.2
-IOP [T2w] 17.1 ± 0.1 17.6 ± 0.1 17.0 ± 0.1 17.9 ± 0.2 18.8 ± 0.3 19.6 ± 0.2 18.0 ± 0.1 19.2 ± 0.1 19.7 ± 0.2
-IOP [PDw] 16.9 ± 0.1 18.3 ± 0.1 18.8 ± 0.1 18.5 ± 0.1 19.8 ± 0.2 20.0 ± 0.3 18.6 ± 0.1 19.8 ± 0.1 20.5 ± 0.2
+HH [T1w] 15.1 ± 0.0 16.5 ± 0.0 19.4 ± 0.0 19.1 ± 0.0 20.0 ± 0.1 20.1 ± 0.1 19.3 ± 0.0 21.0 ± 0.0 21.7 ± 0.0
+HH [T2w] 16.5 ± 0.0 16.9 ± 0.0 16.4 ± 0.0 17.5 ± 0.0 15.6 ± 0.1 18.8 ± 0.0 17.5 ± 0.0 18.5 ± 0.0 18.9 ± 0.0
+HH [PDw] 16.5 ± 0.0 17.8 ± 0.0 17.8 ± 0.0 18.0 ± 0.0 19.2 ± 0.0 18.9 ± 0.1 18.2 ± 0.0 19.3 ± 0.0 19.9 ± 0.0
+IOP [T1w] 14.7 ± 0.0 16.7 ± 0.0 18.9 ± 0.0 18.4 ± 0.0 19.9 ± 0.0 18.5 ± 0.1 18.8 ± 0.0 20.3 ± 0.0 21.0 ± 0.0
+IOP [T2w] 17.1 ± 0.0 17.6 ± 0.0 17.0 ± 0.0 17.9 ± 0.0 18.8 ± 0.0 19.6 ± 0.0 18.0 ± 0.0 19.2 ± 0.0 19.7 ± 0.0
+IOP [PDw] 16.9 ± 0.0 18.3 ± 0.0 18.8 ± 0.0 18.5 ± 0.0 19.8 ± 0.0 20.0 ± 0.0 18.6 ± 0.0 19.8 ± 0.0 20.5 ± 0.0
 5 Discussion
@@
 Our approach faces several limitations.
-First, training on full-resolution 3D volumes is computationally demanding, limiting batch size and reducing the number of negative pairs used during contrastive learning.
+Full-resolution 3D training is costly, so batch size - and thus negative pairs - is limited.
 Second, we rely on a CNN backbone, which may not capture long-range dependencies as effectively as vision transformers or other recent architectures.
 Third, while cross-sequence invariance bolsters model robustness, certain sequence-specific gaps - particularly on T2w images - highlight the need for further improvements.
-Finally, qMRI data are relatively scarce, leading us to pre-train on only 51 subjects; scaling to a larger qMRI dataset or inferring parameter maps from a resource like the UK Biobank (Sudlow et al., 2015) could significantly boost downstream performance.
+Further, qMRI inherently is unable to generate modalities such as SWI, DWI or CT, and therefore may still be liable to domain shifts in the presence of such modalities.
+A notable limitation was pretraining on only 51 subjects, which is relatively small for SSL frameworks.
+Scaling pretraining to larger qMRI datasets or synthetically derived qMRI maps from large databases such as the UK Biobank could further enhance representation robustness.
 5.3 Future Directions
-We plan to explore ViT-based encoders, alternative SSL methods (VICReg, DINO), and larger clinical datasets to further improve domain invariance and push towards routine clinical adoption.
-We also expect multi-view contrastive learning and decoder pretraining (Asiedu et al., 2022) to be valuable directions of future work.
+Future work will test ViT encoders, larger qMRI datasets and alternative SSL objectives such as VICReg and DINO.
+We also expect multi-view contrastive learning and decoder pretraining [2] to be valuable directions of future work.
+By leveraging large public datasets of structural MRI, it may be possible to use existing methods for estimating qMRI such as [6] to generate a large, synthetic dataset to benefit from the scaling effects of self-supervised pre-training.
 5.4 Conclusion
@@
 We believe this framework provides a stepping stone toward truly cross-domain, clinically deployable deep learning models in medical imaging.
-[…]
-LC is supported by the EPSRC-funded UCL Centre for Doctoral Training in Intelligent, Integrated Imaging in Healthcare (i4health) (EP/S021930/1), and the Wellcome Trust (203147/Z/16/Z and 205103/Z/16/Z).
-This research was supported by NVIDIA and used NVIDIA RTX A6000 48GB.
-Appendix A Physics-based Signal Equations
-This appendix details the forward models used to simulate MRI signal intensities from quantitative MRI (qMRI) parameter maps.
-For each voxel, we assume known proton density (PD), longitudinal relaxation rate ($R_{1}$), and transverse relaxation rate ($R_{2}$ or $R_{2}^{*}$), and optionally a magnetic transfer (MT) parameter.
-We consider several common MRI sequences, each governed by its own signal equation:
-Fast Spin-Echo (FSE).
-$S_{\text{FSE}}=B_{1}\cdot\text{PD}\cdot\bigl{(}1-e^{-R_{1}\cdot T_{R}}\bigr{)}\cdot e^{-R_{2}\cdot T_{E}}.$ (1)
-Gradient-Echo (GRE).
-$S_{\text{GRE}}=B_{1}\cdot\text{PD}\cdot\sin(\alpha)\cdot(1-\text{MT})\cdot\tfrac{1-e^{-R_{1}\cdot T_{R}}}{1-\cos(\alpha)(1-\text{MT})\,e^{-R_{1}\cdot T_{R}}}\cdot e^{-R_{2}^{*}\cdot T_{E}}.$ (2)
-Fluid-Attenuated Inversion Recovery (FLAIR).
-$S_{\text{FLAIR}}=B_{1}\cdot\text{PD}\cdot e^{-R_{2}\cdot T_{E}}\cdot\bigl{(}1-2\,e^{-R_{1}\cdot T_{I}}+e^{-R_{1}\cdot T_{R}}\bigr{)}.$ (3)
-Magnetisation-Prepared Rapid Gradient Echo (MPRAGE).
-$S_{\text{MPRAGE}}$ $=\Bigl{|}\sin(\alpha)\,\Bigl{[}\tfrac{1-e^{-R_{1}\cdot T_{R}}}{1-\cos(\alpha)\,e^{-R_{1}\cdot T_{R}}}\,\bigl{(}1-(\cos(\alpha)\,e^{-T_{X}\cdot R_{1}})^{n}\bigr{)}\Bigr{]}\,e^{-T_{D}\cdot R_{1}}$ (4)
-$\quad+\,\bigl{(}1-e^{-T_{D}\cdot R_{1}}\bigr{)}\Bigr{|}.$
-Here, $B_{1}$ represents the receive field strength, $\alpha$ (in radians) is the flip angle, $T_{R}$ the repetition time, $T_{E}$ the echo time, $T_{I}$ the inversion time, $T_{X}$ the excitation repetition time, $T_{D}$ a delay time, and $n$ the number of excitation pulses (all times in seconds unless noted otherwise).
+Acknowledgements.
+LC is supported by the EPSRC CDT in Intelligent, Integrated Imaging in Healthcare (EP/S021930/1) and by the Wellcome Trust (203147/Z/16/Z and 205103/Z/16/Z).
+NVIDIA donated the RTX A6000 48 GB GPU used in this research.
+Appendix 0.A Physics-based Signal Equations
+For each voxel we assume proton density (PD), longitudinal relaxation rate $R_{1}$, transverse relaxation rate $R_{2}$ ($R_{2}^{*}$ for GRE) and, optionally, magnetisation transfer (MT).
+The receive field is denoted $B_{1}$ and the sequence-specific timing parameters are the repetition time $T_{R}$, echo time $T_{E}$, inversion time $T_{I}$, excitation spacing $T_{X}$, delay $T_{D}$ and excitation count $n$.
+Table 4: Forward signal models used for sequence synthesis.
+Sequence Signal equation $S=f(\cdot)$
+Fast Spin-Echo (FSE) $\text{PD}\,B_{1}\bigl{(}1-e^{-R_{1}T_{R}}\bigr{)}\,e^{-R_{2}T_{E}}$
+Gradient-Echo (GRE) $\text{PD}\,B_{1}\sin\!\alpha\,(1-\text{MT})\,\dfrac{1-e^{-R_{1}T_{R}}}{1-\cos\!\alpha\,(1-\text{MT})e^{-R_{1}T_{R}}}\,e^{-R_{2}^{*}T_{E}}$
+FLAIR $\text{PD}\,B_{1}e^{-R_{2}T_{E}}\bigl{(}1-2e^{-R_{1}T_{I}}+e^{-R_{1}T_{R}}\bigr{)}$
+MPRAGE $\text{PD}\,B_{1}\!\Bigl{|}\sin\!\alpha\frac{1-e^{-R_{1}T_{R}}}{1-\cos\!\alpha\,e^{-R_{1}T_{R}}}\bigl{[}1-(\cos\!\alpha\,e^{-T_{X}R_{1}})^{n}\bigr{]}e^{-T_{D}R_{1}}+1-e^{-T_{D}R_{1}}\Bigr{|}$
 Noise Simulation
-To introduce realistic MRI noise, we add Rician noise Gudbjartsson and Patz (1995) to the simulated signals. Specifically,
-$S_{\textrm{noisy}}=\sqrt{(S_{\textrm{MRI}}+\epsilon_{\textrm{r}})^{2}+\epsilon_{\textrm{i}}^{2}}\text{ where }\epsilon_{\textrm{r}}\sim\mathcal{N}(0,\sigma^{2})\text{ and }\epsilon_{\textrm{i}}\sim\mathcal{N}(0,\sigma^{2})$ (5)
-where $S_{\text{MRI}}$ is the noise-free signal, and each $\mathcal{N}(0,\sigma^{2})$ term is a zero-mean Gaussian simulating real and imaginary noise components.
-All simulations were implemented using the NiTorch libraryhttps://github.com/balbasty/nitorch.
-Acquisition Parameters and Data Augmentation
-We randomly sampled MRI sequences and parameters from realistic distributions to produce diverse training data, as summarised below.
-Each sequence was chosen with equal probability to ensure broad coverage of acquisition protocols.
-FLAIR:
-$T_{E}\sim 10^{\mathcal{N}(\log_{10}(0.02),\log_{10}(0.1))}\quad\text{s},\quad T_{R}\sim 10^{\mathcal{U}(\log_{10}(0.001),\log_{10}(5))}\quad\text{s},$
-$T_{I}\sim 10^{\mathcal{U}(\log_{10}(0.001),\log_{10}(3))}\quad\text{s}.$
-FSE:
-$T_{E}\sim 10^{\mathcal{U}(\log_{10}(0.001),\log_{10}(3))}\quad\text{s},\quad T_{R}\sim 10^{\mathcal{U}(\log_{10}(0.001),\log_{10}(3))}\quad\text{s}.$
-MPRAGE:
-$T_{R}\sim\mathcal{N}(23,2.3)\quad\text{s},\quad T_{I}\sim\mathcal{U}(0.6,0.9)\quad\text{s},\quad T_{X}\sim\mathcal{U}(0.004,0.008)\quad\text{s},$
-$T_{E}\sim\mathcal{U}(0.002,0.004)\quad\text{s},\quad\alpha\sim\mathcal{U}(5^{\circ},12^{\circ}).$
-GRE:
-$T_{E}\sim 10^{\mathcal{U}(\log_{10}(0.002),\log_{10}(0.08))}\quad\text{s},\quad T_{R}\sim 10^{\mathcal{U}(\log_{10}(0.005),\log_{10}(5))}\quad\text{s},\quad\alpha\sim\mathcal{U}(5^{\circ},50^{\circ}).$
-All sampled parameters were constrained to physically plausible ranges (e.g.
-taking absolute values as needed).
-This broad variation in sequences and parameter values helps train the model to handle diverse imaging protocols.
-Appendix B Model Architectures
+Rician corruption is applied on-the-fly:
+$S_{\text{noisy}}=\sqrt{\bigl{(}S_{\text{MRI}}+\epsilon_{r}\bigr{)}^{2}+\epsilon_{i}^{2}},\quad\epsilon_{r},\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$
+as in [13]. All signal synthesis relies on NiTorch [^2] https://github.com/balbasty/nitorch.
+Acquisition Parameter Sampling
+Table 5: Sampling ranges for synthetic sequence generation ($\log U$ indicates sampling uniformly in log-space).
+Sequence $T_{E}$ [s] $T_{R}$ [s] Additional parameters
+FLAIR $\log U(0.02,0.10)$ $\log U(0.001,5)$ $T_{I}\sim\log U(0.001,3)$
+FSE $\log U(0.001,3)$ $\log U(0.001,3)$
+MPRAGE $U(0.002,0.004)$ $N(23,2.3)$ $T_{I}\!\sim\!U(0.6,0.9)$,  $T_{X}\!\sim\!U(0.004,0.008)$,  $\alpha\!\sim\!U(5^{\circ},12^{\circ})$
+GRE $\log U(0.002,0.08)$ $\log U(0.005,5)$ $\alpha\!\sim\!U(5^{\circ},50^{\circ})$
+All samples are clamped to physically plausible values (negative draws are reflected).
+Appendix 0.B Model Architectures
 Pre-training Architecture
-The pre-training setup consists of three components:
-  • CNN Encoder:
-    • Input: 3D volume with 1 channel
-    • 5 downsampling blocks with features: (64, 128, 256, 512, 768)
-    • Each block: Two 3D convolutions with kernel size 3
-    • Instance normalization and GELU activation after each convolution
-    • Dropout rate: 0.2
-  • Projector: Two-layer MLP
-    • Input dimension: 768
-    • Hidden layer (768 → 512) with GELU activation
-    • Output projects to contrastive learning space (512 → 128)
-  • Reconstructor: Decoder network
-    • 4 transposed convolution blocks: (768 → 384 → 192 → 96 → 48)
-    • Each upsampling: 2×2×2 stride
-    • Final 1×1×1 convolution to output channels
+The pre-training setup consists of three components, described in Table 6. NT-Xent projection head is
+Table 6: Network modules used during self-supervised pre-training.
+Module Layers / blocks Kernel Output dims Notes
+CNN encoder 5 conv-blocks $3^{3}$ 64 $\rightarrow$ 768 instance-norm, GELU, dropout 0.2
+Projector MLP(768 → 512 → 128) NT-Xent projection head
+Reconstructor 4 transposed conv $2^{3}$ 768 $\rightarrow$ 48 L1 reconstruction branch
 Downstream Task Architectures
@@
     • Output: 1 channel (predicted noise)
-B.1 Training Details
+0.B.1 Training Details
 The models were trained with the following specifications: